Two planes which are 2475 miles apart fly toward each other. Their speed differ by 75mph. If they pass each other in 3 hours, what is the speed of each?
I solved this and got:
1st plane- 400mph
2nd plane- 475mph.
Am I correct?

Answer by Alan3354(69443)

Two planes which are 2475 miles apart fly toward each other. Their speed differ by 75mph. If they pass each other in 3 hours, what is the speed of each?
I solved this and got:
1st plane- 400mph
2nd plane- 475mph.
Am I correct?
================
Close.
The sum of the speeds is 2475/3 = 825 mi/hr
---------
F + S = 825
F - S = 75
------------
2F = 900
F = 450 mi/hr for the Faster plane.
--> 375 for the Slower.